# Exploring Remote Collaboration: The Impact of Avatar Representation on Dyadic Haptic Interactions in Shared Virtual Environments

### v1

@@ V-B Objective measures
Moreover, two studies that examined the impact of visual interaction in remote collaboration tasks found that showing partner manipulation information as cursors on the screen improved the performance of dyad work [46, 47]. The tasks in their studies required estimating the partner’s maneuvers, where visual information about the partner was crucial for task completion. Our plate control task involved coordinating the manipulation of a rigid plate, and the information provided through haptic interaction may have been sufficient to estimate the partner’s actions. If the task had involved moving a table, where hand positioning is critical, avatar representation might significantly impact performance, especially when the partner’s physical information plays a crucial role in task completion.
In terms of effort, hypothesis ${\bf H3}$ was rejected. Dyads did not alter their effort despite changes in visual information. For physically connected dyads, the forces between partners play a crucial role, especially in negotiating and integrating intentions [12, 13]. Groten et al. demonstrated that haptic information provides a valuable channel for integrating intentions between dyads in co-performed haptic tasks. Roche et al. supported this, explaining that when dyad intentions differ, the forces exerted increase, and haptic information is utilized in negotiation. These findings suggest that the presence of an avatar did not influence users’ intentions in performing tasks. This could be because the avatars used in our study were generic and lacked gender, expressions, gaze direction, and physical cues that could significantly influence interactions. Indeed, studies have shown that avatar appearance, such as gender, can influence how much force is perceived as appropriate; for instance, interactions with female avatars might result in gentler forces compared to male avatars [29]. Therefore, using avatars with a higher level of detail could potentially affect haptic interactions.
@@ V-C Generalization
We develop an argument that integrates subjective and objective measures. Although changes in social presence were observed across the four avatar conditions, Performance and Effort remained unchanged.
@@
The first key finding is that for dyads physically connected by haptic interaction, avatar representation does not impact performance or the quality of haptic interactions. This insight could significantly influence the development of physical remote collaboration applications. Implementing tracking systems for such systems may no longer be essential. Typically, developing virtual environments that display avatars requires cameras or sensor-based tracking systems, which are subject to constraints such as location and lighting [3]. Additionally, communication and rendering technologies are needed, which contribute to the cost of system construction in both hardware and software [19]. Thus, our findings could lead to reductions in various costs associated with these systems.
@@
The second key finding is that the enhancement of social presence provided by avatar representation cannot be substituted by haptic interaction alone. This finding is particularly relevant for the entertainment sector, where user experience is crucial. Presenting avatar representations in real time to users in a SVE is sufficient to enhance the experience of emotional closeness, much like audio and video [6]. This suggests that avatars play a starring role in social communication channels. Our findings indicate that, on the contrary, haptic interactions play a supporting role. Consequently, avatar representation might be more important than haptic interaction for enhancing user experience, emphasizing the significance of visual elements in the perception of social presence and interaction quality.
@@
This distinction highlights the critical role avatars play in the user’s emotional and psychological engagement within virtual environments, suggesting a strategic focus on visual representation to enhance social presence effectively, particularly in applications where the social dimension is predominant.
@@ VI-A Level of Detail on avatar representation
In SVEs, avatar expression also has a significant impact on users. A study by Kang et al. reported that a low LoD in facial expressions of avatars can diminish social presence [23]. Conversely, interacting with expressive avatars can enhance social presence [30, 52, 53].
This suggests that varying the appearance and expressions of avatars could extend the discussions presented in this paper. Typically, cameras are used to apply facial expressions to avatars, but integrating them with an HMD presents challenges [54]. To address this, facial tracking devices that can be mounted on HMDs have been developed, allowing expressions to be represented even while wearing an HMD [55, 56]. Recently, Apple has developed a personalized avatar system, known as Persona, which can express facial details using their Vision Pro device [57].
In future studies, we plan to use these advanced devices to expand our investigation and further explore how enhanced avatar realism and expressiveness impact user experience in SVEs. This approach may reveal deeper insights into the behavioral and psychological effects of high-fidelity avatars, potentially guiding the design of more engaging and effective virtual reality experiences.
@@ VI-B The impact on group size
In this paper, participants were limited to dyads, but there is potential to expand the research to groups larger than three people. Generally, as the group size increases, the dynamics of the group become more pronounced, and various psychological effects occur. For instance, larger groups might see an increase in free riders and a decrease in individual proactivity and performance. This decrease can be explained by phenomena such as the Ringelmann effect or the bystander effect [58, 59].
In our investigation, we found only one study that investigated the impact of changes in group size on social presence within an SVE [60]. Their research involved tasks where 2–4 person groups used HMDs to communicate, examining social presence. They found that smaller groups tend to engage more closely with each other, experiencing a higher social presence. In contrast, larger groups might feel a certain distance in their interactions with others, which can affect the sense of social presence.
Given these findings, it is clear that increasing group size could have significant impacts, but actual effects would need to be investigated. Fortunately, our developed system supports collaborative work for up to four people, allowing future studies to explore the effects of group size. This exploration could significantly contribute to understanding how social presence is mediated by group dynamics in virtual environments, potentially influencing the design of virtual reality systems and applications for larger groups.
@@ VI-C Collaboration task
Maroger et al. conducted research on the task of carrying a table in dyads. Carrying a table is a routine and seemingly simple context, but they showed that the strategies for carrying a table vary by dyad and are complex [61]. Sawers et al. conducted research with an interesting task of dancing in dyads [62]. Their study demonstrated the potential for small interaction forces to convey movement goals.
@@ VII CONCLUSION
The purpose of this paper was to investigate how avatar representation affects social presence and collaborative tasks in a Shared Virtual Environment (SVE) with haptic interaction. Four conditions of avatar representation were prepared: both the participant’s and the partner’s avatars are displayed (ALL), only the participant’s avatar is displayed (ALONE), only the partner’s avatar is displayed (OTHER), and no avatars are displayed (NONE). The experiment involved a plate control task with haptic interaction, which provided consistency among conditions. The analysis was carried out through subjective measures of social presence and objective measures of performance and effort. Social presence scored higher in ALL and OTHER conditions and lower in the ALONE and NONE conditions. Therefore, although haptic interaction provided the presence of a partner, the presence of the partner’s avatar was suggested to be important for enhancing social presence. Performance and effort remained consistent regardless of avatar representation. This suggests that avatar representation does not affect haptic interaction. From these results, we present the following contributions: 1) Avatar representation does not impact performance or effort of haptic interaction, therefore avatars can be omitted when the sole purpose is physical collaboration. 2) Visual interaction through avatars plays a leading role in social communication channels, but haptic interaction can only play a supporting role.
We hope that the findings of this paper will aid in the development of physical remote collaboration through SVEs.

### v2

@@ V-B Objective measures
Moreover, two studies that examined the impact of visual interaction in remote collaboration tasks pointed out that showing partner manipulation information as cursors on the screen improved the performance of dyad work [52, 53]. The tasks in their studies required estimating the partner’s maneuvers, where visual information about the partner was crucial for task completion. Our plate control task involved coordinating the manipulation of a rigid plate, and the information provided through haptic interaction may have been sufficient to estimate the partner’s actions. If the task had involved moving a table, where hand positioning is critical, avatar representation might significantly impact performance, especially when the partner’s physical information plays a crucial role in task completion.
In terms of effort, hypothesis ${\bf H3}$ was rejected. Dyads did not alter their effort despite changes in visual information. For physically connected dyads, the forces between partners play a crucial role, especially in negotiating and integrating intentions [12, 13]. Groten et al. demonstrated that haptic information provides a valuable channel for integrating intentions between dyads in co-performed haptic tasks. Roche et al. supported this, explaining that when dyad intentions differ, the forces exerted increase, and haptic information is utilized in negotiation. These findings suggest that the presence of an avatar does not influence users’ intentions in performing tasks. This is because the avatars used in our study were generic and lacked gender, expressions, gaze direction, and physical cues that could significantly influence interactions. However, studies have shown that avatar appearance, such as gender, can influence the degree of force perceived to be appropriate for instance, interactions with female avatars might result in gentler forces compared to male avatars [30]. Therefore, using avatars with a higher level of detail could potentially affect haptic interactions.
@@ V-C Generalization
We developed an argument that integrated subjective and objective measures. Although changes in social presence were observed across the four avatar conditions, performance and effort remained unchanged.
@@
The first key finding is that for dyads physically connected by haptic interaction, avatar representation does not impact performance or the quality of haptic interactions. This insight could significantly influence the development of physical remote collaboration applications. Implementing tracking systems for such systems may not be essential. Typically, developing VEs that display avatars requires cameras or sensor-based tracking systems, which are subject to constraints, such as location and lighting [3]. Additionally, communication and rendering technologies are needed, which contribute to the cost of system construction in both hardware and software [19]. Thus, our findings could cause reductions in various costs associated with these systems.
@@
The second key finding is that the enhancement of social presence provided by avatar representation cannot be substituted by haptic interaction alone. This finding is particularly relevant for the entertainment sector, where user experience is crucial. Presenting avatar representations in real time to users in an SVE is sufficient to enhance the experience of emotional closeness, similar to that of audio and video [6]. Therefore, avatars play a starring role in social communication channels. Our findings indicate that, on the contrary, haptic interactions play a supporting role. Consequently, avatar representation might be more important than haptic interaction for enhancing user experience, emphasizing the significance of visual elements in the perception of social presence and interaction quality.
@@
This distinction highlights the critical role avatars play in the user’s emotional and psychological engagement within VEs, suggesting a strategic focus on visual representation to enhance social presence effectively, particularly in applications where the social dimension is predominant.
Our experiment intentionally employed a controlled and abstract task to isolate the effects of avatar representation and haptic interaction. This design restricted non-verbal or verbal communication, which may not reflect real-world scenarios where participants often rely on multiple communication cues. Although such an experimental setup allowed us to distinct between the contributions of haptic feedback and avatar representation, it also limits the direct applicability of our findings to more naturalistic tasks that combine social and performance demands simultaneously.
For instance, tasks that involve both a strong social component (e.g., negotiation, empathy, or team building) and a high emphasis on performance [54, 55] may benefit from a more nuanced integration of avatars and haptic feedback. In such scenarios, relying on haptic information might not suffice for effective collaboration, while simultaneously, the need for accurate task performance could make the absence of avatars less optimal[51]. Future research should investigate how richer communication channels—including voice, facial expressions, and higher-fidelity avatars—interact with haptic feedback when both social presence and task performance are equally important. Chellali et al. reported that haptic feedback induces changes in verbal communication, leading to an increase in discussions related to force. Such shifts in conversation can, in turn, reciprocally affect how haptic information is shared, and they are likely to occur in highly contextual social scenarios. Meanwhile, Roth et al. compared the time required to reach an agreement in a negotiation role-play—a verbally oriented task—between a real-world setting and a VE using abstract avatars[56]. They found no significant differences, indicating that even abstract avatars can effectively support communication channels. Therefore, investigating interactions that align with the specific context of social tasks is essential for achieving a comfortable remote collaboration.
In summary, while our findings suggest that avatar representation does not affect task performance in a strictly task-centered scenario, and that haptic feedback alone cannot substitute for visual cues in enhancing social presence, these conclusions are based on a controlled context that may not fully capture the complexity of real-world interactions. Researchers and designers should carefully consider whether their target application places greater emphasis on social presence, task performance, or both, and evaluate if additional communication modalities are needed beyond the basic haptic and visual channels examined here.
@@ VI-A Level of Detail on avatar representation
In SVEs, avatar expression has a significant impact on users. A study by Kang et al. reported that a low LoD in facial expressions of avatars can diminish social presence [23]. Conversely, interacting with expressive avatars can enhance social presence [31, 61, 62]. Moreover, using a 3D-scanned avatar was shown to yield a stronger sense of virtual body ownership (VBO) compared to an abstract avatar, underscoring the importance of avatar realism in social interactions [63].
Therefore, varying the appearance and expressions of avatars could extend the discussions presented in this paper. Typically, cameras are used to apply facial expressions to avatars, but integrating them with an HMD is a challenge [64]. Therefore, facial tracking devices that can be mounted on HMDs have been developed, allowing expressions to be represented even while wearing an HMD [65, 66]. Recently, Apple Inc. has developed a personalized avatar system, known as Persona, which can express facial details using their Apple Vision Pro™[67].
In future studies, we plan to use these advanced devices to expand our investigation and further explore how enhanced avatar realism and expressiveness impact user experience in SVEs. This approach may reveal deeper insights into the behavioral and psychological effects of high-fidelity avatars, potentially guiding the design of more engaging and effective VR experiences.
@@ VI-B The impact on group size
In this paper, participants were limited to dyads, but there is potential to expand the research to groups larger than three people. Generally, as the group size increases, the dynamics of the group become more pronounced, and various psychological effects occur. For instance, larger groups might observe an increase in free riders and a decrease in individual proactivity and performance. This decrease can be explained by phenomena, such as the Ringelmann effect or the bystander effect [68, 69].
In our investigation, we found only one study that investigated the impact of changes in group size on social presence within an SVE [70]. Their research involved tasks where 2–4 person groups used HMDs to communicate, examining social presence. They found that smaller groups tend to engage more closely with each other, experiencing a higher social presence. However, larger groups might feel a certain distance in their interactions with others, which can affect the sense of social presence.
Consider these findings, it is clear that increasing group size could have significant impacts, but actual effects would need to be investigated. However, our developed system supports collaborative work for up to four people, allowing future studies to explore the effects of group size. This exploration could significantly contribute to understanding how social presence is mediated by group dynamics in VEs, potentially influencing the design of VR systems and applications for larger groups.
@@ VI-C Collaboration task
Maroger et al. conducted research on the task of carrying a table in dyads. Carrying a table is a routine and seemingly simple context, however, they showed that the strategies for carrying a table vary by dyad and are complex [71]. Sawers et al. conducted research with an interesting task of dancing in dyads [72]. Their study demonstrated the potential for small interaction forces to convey movement goals.
@@ VII CONCLUSION
In this study, we investigated how avatar representation affects social presence and collaborative tasks in a Shared Virtual Environment (SVE) with haptic interaction. Four conditions of avatar representation were prepared: both the participant’s and the partner’s avatars are displayed (ALL), only the participant’s avatar is displayed (ALONE), only the partner’s avatar is displayed (OTHER), and no avatars are displayed (NONE). In all conditions, we used a generic, gender-neutral avatar to eliminate variables, such as facial expressions or gender cues. The experiment involved a plate control task with haptic interaction, which was intentionally designed as an abstract and controlled scenario to ensure consistency among conditions. The analysis was performed through subjective measures of social presence and objective measures of performance and effort. Social presence scored higher in ALL and OTHER conditions and lower in the ALONE and NONE conditions. Therefore, although haptic interaction provided the presence of a partner, the presence of the partner’s avatar was suggested to be important for enhancing social presence. Performance and effort remained consistent regardless of avatar representation, indicating that avatar representation does not affect haptic interaction under these controlled conditions. Consequently, we present the following contributions: 1) Avatar representation does not impact performance or effort of haptic interaction; therefore, avatars can be omitted when the sole purpose is physical collaboration. 2) Visual interaction through avatars plays a leading role in social communication channels, but haptic interaction can only play a supporting role.
We hope that the findings of this paper, although derived from a generic avatar and a controlled plate control task, will aid in the development of physical remote collaboration through SVEs.